Simplify and express each of the following in exponential form:
 3^7 / (3^4 × 3^3 )

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

= 3^7/ (3^4+3) ... [∵(a^m x a^n ) = a^m+n
= 3^7/ 3^7
= 3^7-7 ... [∵a^m÷a^n = a^m-n ]
= 3^0
= 1

Thanks